how many gallons of 20% alcohol solution should be mixed with 5 gallons of a 70% alcohol solution to get a 45% alcohol solution?
Answer by Alan3354(69443)

Same problem, different numbers
=======================================================
how many liters of a 30% salt solution should be mixed with a 80% salt solution to obtain 120 liters of a 60% salt solution?
-----------
t = amount of 30%
e = amount of 80% --- good luck finding 80% salt
----
t + e = 120 (total solution)
30t + 80e = 120*60 ( total salt)
etc